How do i find the x and y coordinate of the vertex, the line of symmetry, and the max/min of f(x) for the problem f(x)=-2^2+2x+7?
Answer by lwsshak3(11628)

How do i find the x and y coordinate of the vertex, the line of symmetry, and the max/min of f(x) for the problem f(x)=-2x^2+2x+7?
**
f(x)=-2x^2+2x+7
complete the square
y=-2(x^2-2x+1/4)+7+1/2
y=-(x-1/2)^2+15/2
this is an equation for a parabola that opens downwards
Its standard form: y=A(x-h)^2+k , (h,k)=(x,y) coordinates of the vertex, A=multiplier which affects width or steepness of the curve
..
vertex:(1/2,15/2)
line or axis of symmetry: x=1/2
maximum: 15/2